# Tropical Gradient Descent

### v1

@@ 5 Discussion
We have introduced the concept of tropical descent as the steepest descent method with respect to the tropical norm. We have demonstrated, both theoretically and experimentally, that it possesses well-behaved properties with respect to a wide class of optimisation problems on the tropical projective torus, including several key statistical problems for the analysis of phylogenetic data. However, our experimental results are stronger than our theoretical results suggest; the optimisation problems studied are location problems, rather than $\triangle$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset. It would therefore be a natural extension of this work to identify how much we can relax the $\triangle$-star-quasi-convexity condition while guaranteeing convergence.
@@
In this paper, we consider tree data with 4 to 8 leaves, and samples of size 10 to 100—at this scale, we have observed (Trop/)SGD to be a significant sacrifice in accuracy with little benefit in computational time. Furthermore, at this scale, the relative merits of SGD versus its tropical counterpart are inconsistent; TropSGD out-performs SGD when solving the linear regression problem, but performs worse when finding centrality statistics. However, for applications to larger datasets, we would expect the stochastic approaches to be more suitable. It would then be important to understand the theoretical differences between SGD and TropSGD in terms of their convergence guarantees in the tropical setting and their comparative performance in practice.
@@
Finally, as shown by Akian et al., (2023), the tropical linear regression problem is polynomial-time equivalent to mean payoff games - two-player perfect information games on a directed graph (Ehrenfeucht & Mycielski, (1979); Gurvich et al., (1988)). This family of problems of particular interest in game theory as they are known to be NP $\cap$ co-NP. While our methods are designed to approximate the minimum rather than compute it exactly, tropical descent may prove to be an efficient tool when solving large-scale mean payoff games.
Acknowledgments
@@
R.T. receives partial funding from a Technical University of Munich–-Imperial College London Joint Academy of Doctoral Studies (JADS) award (2021 cohort, PIs Drton/Monod). A.M. is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1] and a London Mathematical Society Emmy Noether Fellowship.
Appendix A Data Generation
@@
The same 12 datasets are used for each experiment in Section 4. These datasets live in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$; these dimensions correspond to the spaces of trees with 4 and 8 leaves respectively. Branching process datasets are generated using the R package ape (Paradis & Schliep, (2019)) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian. Coalescent data has also been generated using ape and is supported on the space of ultrametric trees, which is a tropical linear space. Gaussian data is generated using NumPy (Harris et al., (2020)), and is supported on the general tropical projective torus.
@@
Appendix B PCA and Logistic Regression
Figure 9: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the PCA objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Figure 10: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the logistic regression objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (Yoshida et al., (2019)) and tropical logistic regression loss function (Aliatimis et al., (2023)). These loss functions take multiple points in $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as argument, and we execute our gradient methods on each argument simultaneously. Figures 9 and 10 show the resulting CDF of our gradient methods for these problems. These problems do not lend themselves well to gradient methods, as the relative log error is rather large for all of our proposed methods. However, for both problems the tropical descent and TropAdamax methods are outperforming the classical methods quite consistently.
@@ Appendix C Learning Rate Tuning
The Adam algorithms—Adam, Adamax and TropAdamax—are run using the default $\beta=(0.9,0.999)$ parameters. We then tune the learning rates used for each method; Gradient Descent, Tropical Descent, SGD, Tropical SGD, Adam, Adamax, and TropAdamax. We perform an initial test of learning rates $\gamma=0.0001,0.0003,0.001,0.003,0.01,0.03,0.1,0.3,1,3,10,30$ over 1000 steps for each algorithm, with 5 random initialisations for each dataset. From this, we narrowed our search to take 18 initializations at 0.5 intervals on the log scale in near optimal regions for each problem, gradient method, and data dimensionality. We then selected the learning rate with minimal mean log relative error in each case; tables 5 - 11 show the selected learning rates.
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 0.135 0.135 0.135 0.0821 0.00409 0.00409 0.0111
100 0.135 0.135 0.0821 0.135 0.00409 0.00409 0.00674
28 10 0.368 0.368 0.368 0.607 0.00674 0.00674 0.0183
100 0.368 0.223 0.368 0.607 0.00674 0.00674 0.0111
@@
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
@@
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 1.65 7.39 0.0821 0.223 0.135 0.223 0.0821
100 1.65 4.48 0.368 0.135 0.0821 0.135 1.00
28 10 1.65 0.368 0.607 0.607 0.00674 0.0111 0.0183
100 2.72 2.72 0.368 0.607 0.0302 0.0821 0.0498
@@
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 0.0821 0.135 0.0302 0.223 0.00674 0.00674 0.0183
100 0.368 0.135 0.607 0.0498 0.00248 0.00409 0.0183
28 10 0.223 0.607 0.223 0.607 0.00409 0.0111 0.0821
100 0.223 0.607 0.223 0.607 0.00409 0.0111 0.0821
@@
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 2.72 4.48 0.607 1.65 0.135 0.135 0.368
100 2.72 4.48 1.00 0.607 0.0821 0.135 0.607
28 10 7.39 4.48 0.368 1.65 0.0183 0.0183 0.223
100 4.48 7.39 0.368 0.607 0.0302 0.0821 0.0821
@@
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 7.39 12.2 1.00 2.72 1.00 1.00 0.607
100 7.39 12.2 1.00 0.368 0.607 1.00 1.00
28 10 20.1 245 7.39 90.0 4.48 4.48 2.72
100 20.1 90.0 4.48 90.0 1.65 7.39 2.72
@@
Data Dim Data Count CD TD SGD TropSGD Adam Adamax TropAdamax
6 10 0.0302 0.0821 0.223 0.0821 0.00248 0.00248 0.0111
100 0.0498 0.0821 0.368 0.0821 0.00248 0.00248 0.0111
28 10 0.368 0.0821 0.0821 0.135 0.0111 0.0302 0.0111
100 0.223 0.607 1.00 12.2 0.0015 0.00409 0.0498
@@ Appendix D Runtimes
Experiments were implemented on a AMD EPYC 7742 node using a single core, 8 GB. Table 12 shows the average time taken per initalisation for each gradient method and optimisation problem.
Stats Prob CD TD SGD TropSGD Adam Adamax TropAdamax
Fermat–Weber 0.18s 0.19s 0.26s 0.26s 0.22s 0.23s 0.26s
Fréchet Mean 0.23s 0.23s 0.30s 0.30s 0.28s 0.28s 0.31s
2-Wasserstein Projection 1.18s 1.18s 1.75s 1.77s 1.23s 1.24s 1.27s
$\infty$-Wasserstein Projection 1.17s 1.17s 1.71s 1.71s 1.21s 1.21s 1.24s
Polytope PCA 0.60s 0.61s 0.78s 0.78s 0.64s 0.64s 0.68s
Logistic Regression 0.45s 0.45s 0.62s 0.62s 0.68s 0.50s 0.53s
Linear Regression 0.22s 0.22s 0.27s 0.28s 0.26s 0.27s 0.30s
Table 12: Average runtimes for 1000 steps of each gradient method when minimising each objective function.
@@ Appendix E Further Figures
Data CD TD SGD TropSGD Adam Adamax TropAdamax
6$\times$10 Branching Data -4.59 -4.56 -4.01 -2.93 -4.58 -4.59 -4.54
6$\times$10 Coalescent Data -4.58 -4.49 -4.06 -1.86 -4.57 -4.58 -4.50
6$\times$10 Gaussian Data -4.59 -4.35 -4.10 -2.63 -4.58 -4.59 -4.32
6$\times$100 Branching Data -4.60 -4.59 -4.36 -3.90 -4.59 -4.59 -4.59
6$\times$100 Coalescent Data -4.57 -4.57 -4.44 -3.14 -4.54 -4.56 -4.55
6$\times$100 Gaussian Data -4.59 -4.59 -4.44 -4.03 -4.59 -4.60 -4.59
28$\times$10 Branching Data -4.57 -3.69 -3.66 -1.99 -4.54 -4.57 -3.79
28$\times$10 Coalescent Data -4.56 -3.72 -4.01 -2.57 -4.52 -4.57 -3.71
28$\times$10 Gaussian Data -4.56 -4.05 -3.71 -2.49 -4.54 -4.57 -4.02
28$\times$100 Branching Data -4.59 -4.58 -3.99 -3.23 -4.58 -4.59 -4.58
28$\times$100 Coalescent Data -4.54 -4.56 -4.14 -2.87 -4.46 -4.57 -4.52
28$\times$100 Gaussian Data -4.59 -4.55 -3.90 -3.01 -4.59 -4.59 -4.58
Table 13: The mean log relative error of each gradient method after 1000 steps for each dataset when minimising the Fréchet mean objective function. The minimal mean errors for each dataset are in bold.
Figure 11: The CDF of the log relative error after 1000 steps for each gradient method across 50 initialisations when minimising the Fréchet mean objective. Each subfigure corresponds to a dataset of different dimensionality (columns) and distribution (rows).

### v2

@@ 5 Conclusion and Discussion
Our theoretical results outline the relative strengths of tropical descent over classical descent for tropical location problems, but only provide global solvability for 1-sample problems. In practice, our experimental results go further; the optimization problems studied are location problems, rather than $\triangle$-star-quasi-convex problems, and yet tropical descent appears to consistently converge to a global minimum, not just the tropical convex hull of our dataset. It would therefore be a natural extension of this work to identify how much we can relax the $\triangle$-star-quasi-convexity condition while guaranteeing convergence.
@@
In this paper we have considered tree data with 4 to 8 leaves, and samples of size 10 to 100—at this scale, we have observed (T/)SGD to be a significant sacrifice in accuracy with little benefit in computational time. Furthermore, at this scale, the relative merits of SGD versus its tropical counterpart are inconsistent; TSGD out-performs SGD when solving the linear regression problem, but performs worse when finding centrality statistics. In a similar vein, initial investigations suggest that the classical stochastic algorithm for the computation of Fréchet means—Sturm’s algorithm (Sturm,, 2003)—performs poorly for the computation of tropical Fréchet means due to the inconsistent curvature of the tropical projective torus (Améndola & Monod,, 2023). However, for applications to larger datasets and more complex objective functions, stochastic gradient methods may become necessary. It would then be important to understand the theoretical differences between SGD and TSGD in terms of their convergence guarantees in the tropical setting and their comparative performance at different scales of data dimension.
The recent identification of tropical rational functions and ReLU neural networks (Zhang et al.,, 2018) has inspired new avenues of research in tropical applications, into the tropical geometry of neural networks (Brandenburg et al.,, 2024; Yoshida et al.,, 2024). It is therefore natural to consider the possible extensions of this work to neural network training; can we identify local tropical quasi-convexity in the loss functions of ReLU neural networks, and how can we extend tropical steepest descent to the space of tropical rational functions?
@@
Finally, as shown by (Akian et al.,, 2023), the tropical linear regression problem is polynomial-time equivalent to mean payoff games—two-player perfect information games on a directed graph (Ehrenfeucht & Mycielski,, 1979; Gurvich et al.,, 1988). This family of problems is of particular interest in game theory as they are known to be NP $\cap$ co-NP. In our numerical experiments, tropical gradient methods demonstrate a particularly strong performance in solving tropical linear regression, and while our methods are designed to approximate the minimum rather than compute it exactly, tropical descent may prove to be an efficient tool when solving large-scale mean payoff games.
Acknowledgements
@@
R.T. receives partial funding from a Technical University of Munich–Imperial College London Joint Academy of Doctoral Studies (JADS) award (2021 cohort, PIs Drton/Monod). A.M. is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1] and a London Mathematical Society Emmy Noether Fellowship [EN-2223-01].
Declarations
The authors have no relevant financial or non-financial interests to disclose. The authors have no competing interests to declare that are relevant to the content of this article. All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript. The authors have no financial or proprietary interests in any material discussed in this article.
Appendix A PCA and Logistic Regression
Figure 9: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the PCA objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
Figure 10: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the logistic regression objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
As well as the location problems outlined in this paper, we performed the same numerical experiments with the tropical polytope PCA loss function (Yoshida et al.,, 2019) and tropical logistic regression loss function (Aliatimis et al.,, 2024). These loss functions take multiple points in $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as argument, and we execute our gradient methods on each argument simultaneously. Figures 9 and 10 show the resulting CDF of our gradient methods for these problems. These problems do not lend themselves well to gradient methods, as the relative log error is rather large for all of our proposed methods. However, for both problems the tropical descent and TrAdamax methods are outperforming the classical methods quite consistently.
Appendix B Methodology
@@
The same 12 datasets are used for each experiment in Section 4. These datasets live in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ or $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$; these dimensions correspond to the spaces of trees with 4 and 8 leaves respectively. Branching process datasets are generated using the R package ape (Paradis & Schliep,, 2019) with exponentially distributed branch lengths—this produces data supported on the space of metric trees, i.e., the tropical Grassmannian. Coalescent data has also been generated using ape and is supported on the space of ultrametric trees, which is a tropical linear space. Gaussian data is generated using NumPy (Harris et al.,, 2020), and is supported on the general tropical projective torus.
@@
We use PyTorch (Paszke et al.,, 2019) for the computation and differentiation of each objective function. Max-tropical descent directions are used for tropical gradient methods; we therefore optimize the min-tropical location problems—Wasserstein projections and tropical linear regression—with respect to $-\mathbf{t}$.
@@ Appendix C Learning Rate Tuning
The Adam algorithms — Adam, Adamax and TrAdamax — are run using the default $\beta=(0.9,0.999)$ parameters. We then tune the learning rates used for each method; Gradient Descent, Tropical Descent, SGD, Tropical SGD, Adam, Adamax, and TrAdamax. We perform an initial test of learning rates $\gamma$ on a logarithmic scale from 0.0001 to 30 over 1000 steps for each algorithm, with 5 random initializations for each dataset. From this, we narrowed our search to take 18 initializations at 0.5 intervals on the log scale in near optimal regions for each problem, gradient method, and data dimensionality. We then selected the learning rate with minimal mean log relative error in each case; Tables 3, 4, 5, 6, 7, 8 and 9 show the selected learning rates.
@@
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
@@
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
6 10 0.135 0.135 0.135 0.135 0.00674 0.00674 0.00674
100 0.135 0.135 0.135 0.135 0.00674 0.00248 0.00674
28 10 0.368 0.368 0.368 1 0.00674 0.00674 0.0183
100 0.368 0.135 0.368 1 0.00674 0.00674 0.0183
@@
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
6 10 2.72 1 1 0.135 0.135 0.368 0.368
100 2.72 2.72 0.368 0.135 0.135 0.135 0.135
28 10 0.368 0.368 0.368 1 0.00674 0.0183 0.0183
100 2.72 2.72 0.368 1 0.0498 0.135 0.0498
@@
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
6 10 0.0498 0.135 0.0498 0.368 0.00674 0.00674 0.0183
100 0.0498 0.135 0.135 2.72 0.00674 0.00674 0.0183
28 10 0.368 1 0.368 1 0.00674 0.0183 0.0498
100 0.368 1 0.368 1 0.00674 0.0183 0.0498
@@
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
6 10 2.72 2.72 1 1 0.135 0.135 0.368
100 2.72 7.39 1 1 0.135 0.135 1
28 10 7.39 7.39 0.368 1 0.0183 0.0183 0.135
100 2.72 7.39 0.368 1 0.0498 0.0498 0.0498
@@
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
6 10 7.39 2.72 1 2.72 0.368 1 0.0498
100 1 0.368 1 2.72 0.00248 0.00248 0.00248
28 10 1 0.368 1 2.72 0.00248 0.00248 0.00674
100 1 0.368 1 7.39 0.00248 0.00248 0.00674
@@
Data Dim Data Count CD TD SGD TSGD Adam Adamax TrAdamax
6 10 0.368 0.135 0.368 0.135 0.00248 0.0183 0.00674
100 0.135 0.135 0.368 0.0498 0.00248 0.00248 0.00674
28 10 0.368 0.0498 0.0498 0.135 0.00248 0.00248 0.0183
100 1 0.368 1 7.39 0.0498 0.0498 0.0498
@@ Appendix D Runtimes
Experiments were implemented on a AMD EPYC 7742 node using a single core, 8 GB. Table 10 shows the average time taken per initialization for each gradient method and optimization problem.
Table 10: Average runtimes for 1000 steps of each gradient method when minimizing each objective function.
Stats Prob CD TD SGD TSGD Adam Adamax TrAdamax
Fermat–Weber 0.480 0.520 0.435 0.499 0.391 0.460 0.464
Fréchet Mean 0.347 0.363 0.301 0.342 0.294 0.422 0.370
2-Wasserstein Projection 2.739 2.134 2.392 2.209 2.341 2.246 2.272
$\infty$-Wasserstein Projection 2.947 3.049 2.772 2.565 2.501 3.162 2.807
Polytope PCA 1.117 1.081 1.014 1.052 1.037 1.174 1.029
Logistic Regression 0.987 1.015 0.962 0.934 0.943 1.084 0.943
Linear Regression 0.494 0.499 0.386 0.420 0.386 0.537 0.433
@@ E.1 Fréchet Means
Table 11: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the Fréchet mean objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TSGD Adam Adamax TrAdamax
6$\times$10 Branching Data -4.59 -4.54 -3.98 -2.99 -4.58 -4.59 -4.53
6$\times$10 Coalescent Data -4.58 -4.47 -4.07 -1.82 -4.57 -4.58 -4.45
6$\times$10 Gaussian Data -4.59 -4.33 -4.09 -2.63 -4.58 -4.59 -4.35
6$\times$100 Branching Data -4.59 -4.59 -4.35 -4.03 -4.59 -4.58 -4.59
6$\times$100 Coalescent Data -4.57 -4.56 -4.36 -3.29 -4.55 -4.56 -4.57
6$\times$100 Gaussian Data -4.59 -4.60 -4.39 -4.12 -4.59 -4.58 -4.59
28$\times$10 Branching Data -4.58 -3.73 -3.78 -1.98 -4.55 -4.58 -3.83
28$\times$10 Coalescent Data -4.56 -3.67 -3.98 -2.46 -4.53 -4.57 -3.70
28$\times$10 Gaussian Data -4.55 -4.07 -3.91 -2.41 -4.55 -4.57 -4.04
28$\times$100 Branching Data -4.59 -4.58 -4.03 -2.98 -4.58 -4.59 -4.57
28$\times$100 Coalescent Data -4.54 -4.56 -4.08 -2.66 -4.47 -4.57 -4.46
28$\times$100 Gaussian Data -4.59 -4.56 -4.08 -2.78 -4.59 -4.59 -4.58
Figure 11: The CDF of the log relative error after 1000 steps for each gradient method across 50 initializations when minimizing the Fréchet mean objective. Each sub-figure corresponds to a dataset of different dimensionality (columns) and distribution (rows).
E.2 Wasserstein Projections
From Tables 12 and 13, we see that when solving the 2-Wasserstein projection problem, tropical methods are optimal in all cases other than $28\times 10$ coalescent data. For the $\infty$-Wasserstein projection problem in small dimensions, tropical descent outperforms other methods by some margin, while the comparative performances are less consistent for the $\infty$-Wasserstein projection problem in high dimensions.
Table 12: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the $2$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TSGD Adam Adamax TrAdamax
6$\times$10 Branching Data -3.02 -4.09 -2.16 -2.11 -2.95 -3.00 -3.89
6$\times$10 Coalescent Data -3.71 -3.92 -2.44 -3.13 -3.78 -3.68 -4.12
6$\times$10 Gaussian Data -3.59 -3.83 -2.87 -2.75 -3.61 -3.57 -3.89
6$\times$100 Branching Data -3.44 -4.24 -2.91 -1.60 -3.79 -3.79 -4.02
6$\times$100 Coalescent Data -4.26 -4.55 -3.27 -2.58 -4.39 -4.45 -4.50
6$\times$100 Gaussian Data -3.69 -4.17 -3.33 -2.73 -3.78 -3.78 -4.18
28$\times$10 Branching Data -3.39 -4.31 -3.16 -2.03 -3.53 -3.55 -4.32
28$\times$10 Coalescent Data -4.11 -3.79 -3.39 -2.42 -4.17 -4.12 -3.80
28$\times$10 Gaussian Data -3.43 -3.91 -3.21 -2.15 -3.50 -3.44 -3.83
28$\times$100 Branching Data -3.61 -4.55 -3.37 -2.65 -3.85 -3.85 -4.51
28$\times$100 Coalescent Data -4.11 -4.57 -3.89 -3.32 -4.21 -4.19 -4.56
28$\times$100 Gaussian Data -3.66 -4.52 -3.41 -2.61 -3.82 -3.80 -4.49
Table 13: The mean log relative error of each gradient method after 1000 steps for each dataset when minimizing the $\infty$-Wasserstein projection objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TSGD Adam Adamax TrAdamax
6$\times$10 Branching Data -1.63 -4.32 -0.57 -0.91 -1.68 -1.69 -4.32
6$\times$10 Coalescent Data -4.00 -4.47 -0.99 -1.43 -4.03 -4.04 -4.46
6$\times$10 Gaussian Data -3.77 -4.42 -1.76 -3.33 -3.76 -3.75 -4.40
6$\times$100 Branching Data -3.35 -4.48 -1.94 -0.95 -3.34 -3.36 -4.47
6$\times$100 Coalescent Data -3.85 -4.51 -2.06 -2.21 -3.84 -3.85 -4.51
6$\times$100 Gaussian Data -3.63 -4.40 -1.61 -1.22 -3.70 -3.63 -4.36
28$\times$10 Branching Data -3.68 -3.53 -2.22 -1.17 -4.11 -3.93 -3.17
28$\times$10 Coalescent Data -4.22 -3.95 -2.76 -1.92 -4.42 -4.36 -3.94
28$\times$10 Gaussian Data -4.02 -3.14 -2.13 -0.95 -4.43 -4.30 -2.85
28$\times$100 Branching Data -3.46 -3.58 -1.37 -1.19 -3.68 -3.59 -3.39
28$\times$100 Coalescent Data -4.41 -3.92 -2.09 -2.00 -4.54 -4.51 -3.81
28$\times$100 Gaussian Data -2.98 -3.30 -1.64 -1.61 -3.07 -3.04 -2.80